OpenStudy (anonymous):

which radical expression is equivalent to 5^3/2

OpenStudy (solomonzelman):

\[\huge\color{blue}{ a^{(b/c)}=\sqrt[c]{a^b} } \]

OpenStudy (solomonzelman):

\[\huge\color{red}{ 5^{(3/2)}=~?~ } \]

OpenStudy (solomonzelman):

I see you get it, becuase you gave me a medal, no problem, glad I helped.

OpenStudy (anonymous):

5 3/2